# MDAR: A Multi-scene Dynamic Audio Reasoning Benchmark

### v1

@@ 4.1 Setting
We evaluated the performance of both cascaded and non-cascaded models on MDAR. The non-cascaded models include large audio-language models, large audio-reasoning models, and full-modal language models. The large audio-language models consist of Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), SALAMONN (Tang et al., 2024), Audio Flamingo 2 (Ghosh et al., 2025), Audio Flamingo 3 Chat (Goel et al., 2025), MiDashengLM (Dinkel et al., 2025), etc. The large audio-reasoning models include Audio-Reasoner (Xie et al., 2025), etc. The full-modal language models include Omni-R1 (Zhong et al., 2025), Qwen2.5-Omni (Xu et al., 2025a), etc. The cascaded models based on captions include GPT-4o Audio+ Qwen2-Audio-Instruct, Qwen2-Audio-Instruct + Llama-3-Instruct, etc. Of course, the models selected for evaluation vary depending on the question type. This is partly because some models lack sufficient Chinese capabilities or do not support Chinese audio understanding, and partly because some existing models do not accept multimodal inputs with multiple voices. Note that since our benchmark is automatically synthesized by Gemini, the results of Gemini will not be included in the comparison. Regarding the selected and unselected models as well as detailed implementation details, please refer to the Appendix G.
4.2 Results and Discussion
Tables 3.3,  2, and 3 show the results of mainstream models on all tasks of MDAR-main, MDAR-open, and MDAR-multi. Results for the subtasks of MDAR-multi are provided in the Appendix H.
Our Benchmark is Challenging. From the three tables, we can observe that the models’ performance on the bench is as follows: the best-performing model in the multiple-choice section, Qwen2.5-Omni, achieved only 76.67%. The highest score in open-ended questions was only 7.46, and the highest score among open-source models was only 6.58. In the benchmark, the majority of models scored low, with many models having single-digit accuracy rates. This indicates that MDAR is highly difficult and extremely challenging. Moreover, there is a significant difference in performance among different models. In the multiple-choice section, models such as Audio Flamingo 3 Chat and Kimi-Audio-Instruct did not exceed 20%. The scores in multiple-choice and open-ended questions varied by more than 340%, demonstrating that MDAR has good discriminability.
The Gap Between Open-Source and Closed-Source Models Remains Significant. According to Figure 5, we observed that only performance of Qwen2.5-Omni in the multiple-choice benchmark exceeded that of GPT-4o Audio. The performance of other open-source models still could not surpass that of the closed-source model, and there was a significant gap. This phenomenon was even more pronounced in the open-ended question-answering benchmark and the multiple-choice benchmark. Score of GPT-4o Audio in the open-ended question-answering benchmark was 13.37% higher than the best open-source model, and its metrics in the multiple-choice benchmark exceeded those of the best open-source model by 92.78%. This indicates that open-source models still have significant shortcomings in generating answers and handling multiple inputs.
@@ 4.2 Results and Discussion
Table 3: Results of models on the MDAR-multi benchmark for audio language models are presented. Detailed results for each category can be found in the Appendix H The best results are highlighted in bold and the second-best is underlined. The explanations for the four metrics are provided in $3.$
@@
Figure 5: (a) Left: The gap between open-source and closed-source models on three testing benchmarks. The closed-source models significantly outperform the open-source models. (b) Right: Comparison of different scenarios with the same caption model and the same reasoning model in the cascaded models.
@@
Analysis of Perception and Reasoning Capabilities of Cascaded caption Models. As shown in Figure 5, the experiments focused on the single-choice benchmark using GPT-4o Audio (the best closed-source model) and Qwen2-Audio-Instruct (the baseline open-source model). These two models were used as caption models and combined with other models for reasoning. The results in the latter half of Table 3.3 show that: (i) Under the same caption conditions, combinations of GPT-4o Audio with other models performed worse than GPT-4o Audio alone, indicating its strong perception capabilities. (ii) Combinations of Qwen2-Audio-Instruct with other models outperformed Qwen2-Audio-Instruct alone, suggesting its weaker perception capabilities. (iii) Among the same reasoning models, the combination of GPT-4o Audio and Qwen2.5-Omni outperformed the combination of Qwen2-Audio-Instruct and Qwen2.5-Omni, proving that GPT-4o Audio has stronger perception capabilities. DeepSeek-V3, when used as a reasoning model, achieved results that far exceeded those obtained by combining the same caption model with other reasoning models, and was even comparable to the best non-cascaded models, demonstrating its significant reasoning potential.
Figure 6: Distribution of error types in the responses of four models. Reasoning errors remain the primary type of errors in the responses of LALMs, followed by perception errors.
Figure 7: When Gaussian noise was used to replace the audio on MDAR to test GPT-4o Audio, Qwen2-Audio-Instruct, Qwen2.5-Omni, and Qwen2-Audio-Instruct + Qwen2.5-Omni, the performance of all models experienced a significant drop.
Exploring the Limitations of Perception and Reasoning Capabilities. To explore whether the limitations of current models lie in perception or reasoning, and to identify the capabilities, deficiencies, and potential future improvement directions of mainstream LALMs, we conducted two experiments:
@@
Gaussian Noise Replacement Experiment: Taking four different types of models as representatives, we replaced the audio with Gaussian white noise of the same length and input it into the audio-language models. As shown in Figure 7, the accuracy of the models dropped significantly when they received white noise, approaching random guessing. This indicates that the models received audio information. However, the results also show that even with noisy audio, the models were still able to derive some answers from the textual questions.
Error Analysis: We used Gemini-2.5-flash to categorize the incorrect answers from four models—GPT-4o Audio, Qwen2-Audio-Instruct, Qwen2.5-Omni, and Qwen2-Audio-Instruct + Qwen2.5-Omni—into four types of errors: perception errors, reasoning errors, knowledge gaps, and formatting errors. The results in Firgure 6 revealed that reasoning errors were the primary source of mistakes for most models. For example, the reasoning error rate of GPT-4o Audio was as high as 86.68%, while Qwen2-Audio-Instruct had the highest perception error rate, indicating that its perception capabilities need further enhancement. Knowledge gap errors were relatively rare in models with more than 7B parameters, likely because fundamental audio segments have been thoroughly learned during large-scale pretraining. Regarding formatting errors, different models exhibited varying output preferences in our experiments. For instance, Omni-R1 tended to add line breaks in its answers, and some models could only partially restate the options instead of providing complete answers. These issues reflect that audio models still have room for improvement in their ability to follow instructions. Detailed analysis and case study are provided in the Appendix I and Appendix J.
@@ 5 Conclusion
We introduce MDAR, a benchmark for evaluating models on challenging, multi-scene, dynamically evolving audio reasoning tasks. MDAR comprises 3,000 carefully curated question–answer pairs linked to diverse audio clips, covering five categories of complex reasoning across three question types. We evaluated 26 audio models, revealing that MDAR poses substantial challenges across question formats and evaluation methods. Furthermore, we analyzed gaps between open-source and closed-source models, cascaded caption models’ perception and reasoning capabilities, and overall limitations in model audio understanding. These findings underscore the challenges posed by MDAR and its value as a benchmark for advancing audio reasoning research.
6 Ethical Statement
This study strictly adheres to the ethical norms of academic research. All models and data used are collected in accordance with the principle of fair use, and do not contain identifiable personal information. During the model evaluation process, no direct harm was inflicted on any specific group, nor was any discriminatory or offensive content introduced.
7 Reproducibility Statement
In order to ensure the reproducibility of the research, we have detailed the data sources, model information, hyperparameter configurations, evaluation strategies, and evaluation metrics in the paper. All experiments were conducted in the specified hardware and software environment. We open source the benchmark and evaluation scripts, and the complete data will be released upon acceptance of the paper, ensuring that other researchers can replicate the experimental results under the same conditions and further carry out research based on this.
Appendix A Usage of LLM
During the process of writing the paper, we only used the Large Language Model (LLM) to polish some of the text in order to enhance the fluency and accuracy of the language expression. All research content, viewpoints, and data are original. The LLM was used only as an auxiliary tool and did not participate in the creation of the core content.
Appendix B Task
B.1 Task Introduction
@@
B.2 Task Difficulty
Based on the experimental data in Table 3.3, Table 2, Table 3, and Table 5, it can be found that: among the core task types of single-choice questions and open-ended questions, the scene understanding task is the most difficult, while the anomaly detection and event reasoning tasks also present certain challenges. Further analysis by question type shows that: in the context of open-ended questions, temporal reasoning is the most difficult task type, followed by the anomaly detection task in terms of difficulty; in the context of multiple-choice questions, the difficulty coefficient of the continuous scene reasoning task reaches the highest level. This result indicates that the current model still has significant limitations in its reasoning ability when dealing with dynamically switching continuous scenes, which is a key direction that requires focused breakthroughs in subsequent optimizations.
Appendix C Data Source
The MDAR Metadata Movie Collection includes 500 carefully selected Chinese films, covering a wide range of genres and themes. It is designed to construct complex and dynamic scenarios, so as to meet the requirements of multi-threaded narratives and high production standards. Below is a detailed statistical breakdown of the collection. The 500 films in the MDAR Metadata Movie Collection are categorized by common film genres and themes, with the statistical data presented in the Table 4.
@@ Appendix C Data Source
The MDAR Metadata Movie Collection has the following characteristics:
@@
This data provides a solid foundation for the construction of the MDAR benchmark and also offers abundant resources for relevant research.
@@ Appendix D Prompt of Data construction pipeline
Prompt of Data Generation in MDAR-main System: You are a helpful AI assistant, designed to provide useful QA pair to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the provided video, audio, audio translation, audio description, and video description, create three questions about the audio. These should be three audio reasoning questions with options and answers. The audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. The questions must meet the following criteria: 1.They should be reasoning questions related to the audio. 2.The category of the questions should fall into one of the following: temporal reasoning task, social intent reasoning task, anomaly detection and safety task, and specify which category it belongs to. 3.The posed questions should include a thought chain explanation. 4.The questions cannot be answered solely by the text of the title, they must require reasoning based on both the question and the audio. 5.The questions should introduce some additional reasoning information that does not exist in the audio to increase difficulty. 6.Various information from the audio should not appear directly in the question text; just pose the question. 7.The answers provided should follow this format: Question:….. Question Category:….. Question Options:…… Question Answer:…… Assistant: {response}
@@
Prompt of Data generation in MDAR-multi System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the two provided relevant audio segments along with their respective translations, descriptions, and video descriptions, formulate multiple-choice audio reasoning questions-options-answers-reasoning chains for the audio pair. Audio translation 1 is [audio translation 1], audio description 1 is [audio description 1], video description 1 is [video description 1], audio translation 2 is [audio translation 2], audio description 2 is [audio description 2], video description 2 is [video description 2]. The questions must meet the following criteria: 1. The question must relate to audio pairs and cannot include any information or terms related to the video. 2. The category of the question should belong to one of the following categories, and the category must be specified: continuous scene reasoning task, multi-character interaction reasoning task, plot development reasoning task, anomaly detection and security task, time reasoning task. 3. The question should not have an answer that can be derived solely from the text title; it should require reasoning that combines the question and the audio. 4. The question should involve multi-step reasoning and be challenging, with options of moderate length for the answers. 5. The multiple-choice answers can independently reason for each audio’s description but must include at least one reasoning question that is related to both audio segments simultaneously. 6. Your provided answers must conform to the following format: Question:….. Question Category: …… Question Options: …… Question Answer: …… Question Reasoning Chain: …… Assistant: {response}
@@
Prompt of Data generation in MDAR-open System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Propose an open-ended question for audio based on the given video, audio, audio translation, audio description, and video description. The requirement is an open-ended reference answer thought chain pair, where the audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. At the same time, the questions raised must meet the following requirements: 1. The requirements are only open-ended reasoning questions related to audio, and the video and provided text are only for reference in question setting and cannot be used for question setting (already questions, answers) and thought chain analysis. The questions can only be answered by understanding and reasoning about the audio. 2. The category of the question belongs to one of the following categories, and indicate which task it belongs to: time reasoning task, scene understanding task, character relationship and social reasoning task, event reasoning task, anomaly detection and security task. 3. The answer to the question cannot be found solely from the text question, it needs to be combined with the question and audio for reasoning. 4. The answer to an open-ended question is not unique, just provide a reference answer. The answer you provide should meet the following format: Question: …. Question Category: …. Reference answer: …. Assistant: {response}

### v2

@@ 3.2 Data construction pipeline
3.3 Evaluation metrics
In this section, we introduce the evaluation methods for CMDAR. In order to maintain fairness and consistency, we adopt different evaluation methods for different question types.
CMDAR-main. We use accuracy as the evaluation metric. Since LALMs tend to response option content, we adopt the same response processing pipeline as MMAU, using robust regular expressions to match the answer strings with the given options and calculate the accuracy. For each sample, $acc=1$ if string_match(response,answer)=Truestring\_match(response,answer)=True, otherwise, $acc=0$.
CMDAR-open. We employ a state-of-the-art LLM as an automated evaluator. The evaluator rates the model’s answer on a scale of 0 to 10 based on the given scoring prompt. These prompts consider the usefulness, relevance, accuracy, and comprehensiveness of the answer. Detailed prompt design and examples are provided in the Appendix C. We reduce the randomness of the scoring by conducting multiple evaluations and swapping the positions of the answers.
CMDAR-multi. We use the following four metrics to measure model performance inspired by SATA-BENCH (Xu et al., 2025b), i) Exact Match (EM): For each sample, if the model’s predicted answer is as same as the correct answer, the accuracy is counted as 1; otherwise, it is counted as 0. The final EM is calculated as the average accuracy.
$EM=\frac{1}{N}\sum_{i=1}^{N}1\ (if\ P_{i}=T_{i})$ (1)
ii) Jaccard Index (JI): For each sample, we calculate the intersection of the predicted and true answers divided by their union, and then take the average of all samples to measure the degree of overlap.
iii) Mean Average Precision (Precision): For each sample, we calculate the proportion of correctly predicted answers out of all predicted answers, and then take the average of all samples.
iv) Mean Average Recall (Recall): For each sample, we calculate the proportion of correctly predicted answers out of all true answers, and then take the average of all samples.
Figure 4: Left: Comparison of different combinations with the same caption model or the same reasoning model in the cascaded models. Middle: When Gaussian noise is used to replace the audio on CMDAR-main to test four models, the performance of all models experience a significant drop. Right: Distribution of error types
Table 3: Pearson correlation coefficient and t-test results of the correlation between LLM and human ratings.
Sample Qwen2-Audio-Instruct Qwen2.5-Omni GPT4o-Audio
Pearson 50 0.95312 0.9268 0.9714
Sample Qwen2-Audio-Instruct v.s. Qwen2.5-Omni Qwen2.5-Omni v.s. GPT4o-Audio GPT4o-Audio v.s. Qwen2-Audio-Instruct
t-test 500 T: 6.421, P: 0.000 T: -11.866, P: 0.000 T: -8.389, P: 0.000
4 Experiments
@@ 4.1 Setting
We evaluate the performance of both cascaded and non-cascaded models on CMDAR. The non-cascaded models include large audio-language models, large audio-reasoning models, and full-modal language models. The large audio-language models consist of Qwen-Audio-Chat (Chu et al., 2023), Qwen2-Audio-Instruct (Chu et al., 2024), SALAMONN (Tang et al., 2024), Audio Flamingo 2 (Ghosh et al., 2025), Audio Flamingo 3 Chat (Goel et al., 2025), MiDashengLM (Dinkel et al., 2025), etc. The large audio-reasoning models include Audio-Reasoner (Xie et al., 2025), etc. The full-modal language models include Omni-R1 (Zhong et al., 2025), Qwen2.5-Omni (Xu et al., 2025a), etc. The cascaded models based on captions include GPT-4o Audio + Qwen2-Audio-Instruct, Qwen2-Audio-Instruct + Llama-3-Instruct, etc. Considering the scarcity of Chinese audio reasoning data, some models lack of ability of Chinese reasoning or support for Chinese audio understanding. We annotate in the Table 1 whether they were trained using Chinese corpus to ensure fairness in evaluation. Please note that as our benchmark tests are automatically synthesized by Gemini, the results of Gemini related models will not be included. For the selection of models and detailed information, refer to the Appendix F.
4.2 Results
In this section, we preliminarily demonstrate the performance of various models in CMDAR, which is shown in Tables 1,  2 and 4.
Our Benchmark is Challenging. We incorporate the results of human evaluation and observe that the best model Qwen2.5-Omni-7B in CMDAR-main only achieve an accuracy of 76.67%. The highest score in CMDAR-open is only 7.46, which open-source model is also only 6.58. In CMDAR-multi, the scores of most models are very low, even less than 10%, indicating that CMDAR is difficult and challenging. In addition, the performance differences between different models are also significant, indicating that CMDAR has good discriminative ability.
To ensure the correctness and rationality of using LLM for open scoring, we randomly select 10% of samples and manually score the answers of the Qwen2-Audio-Instruct-7B, Qwen2.5-Omni-7B, and GPT-4o Audio. The manual scoring range is 0-5 points, with 5 points indicating complete correctness and clear expression. We calculate Pearson correlation coefficients, and all three coefficients are greater than 0.9, indicating that our human evaluation is highly correlated and reliable with LLM evaluation. Meanwhile, we conduct paired sample t-tests on these three models based on CMDAR-open, and all p-values are less than 0.05, indicating significant differences in evaluation scores among different models. Table 3 present these two results.
The Gap Between Open-Source and Closed-Source Models Remains Significant. we observe that most open-source models still have significant differences compared to the closed-source model GPT-4o Audio in CMDAR-main. This phenomenon is more pronounced in CMDAR-open and CMDAR-multi. The score of GPT-4o Audio in CMDAR-open is 13.37% higher than the best open-source model, and its indicators in CMDAR-multi are 92.78% higher than the best open-source model. These indicates that open-source models still have significant limitations in exploring perception and reasoning abilities.
@@ 4.2 Results
Table 4: Results of audio language models on the CMDAR-multi benchmark are presented. Detailed results for each category can be found in the Appendix G. The best results are highlighted in bold and the second-best is underlined. The explanations for the four metrics are provided in $3.$
@@
4.3 Discussion
In order to explore deeper limitations and identify the abilities shortcomings, and potential improvement directions of LALMs, we design three experiments for further analysis. Finally, we propose suggestions for the future development of LALMs.
@@ 4.3 Discussion
Analysis of Perception and Reasoning Capabilities of Cascaded caption Models. As shown in Figure 4, GPT-4o Audio (best closed-source model) and Qwen2-Audio-Instruct (baseline open-source model) are used as caption models, combined with other reasoning models for CMDAR-main. It can observe that i) Under the same caption, the combination of GPT-4o Audio and other models performs worse than using GPT-4o Audio alone, indicating that GPT-4o Audio has strong perceptual ability. ii) The combination of Qwen2-Audio-Instruct and other models is superior to using Qwen2-Audio-Instruct alone, indicating that the perceptual ability of Qwen2-Audio-Instruct is weaker. iii) When DeepSeek-V3 is used as a reasoning model, its results outperform those combined the same caption model with other reasoning models, and can be comparable to the best non-cascaded models, demonstrating its significant reasoning potential.
@@
Gaussian Noise Replacement Experiment: Taking four different types of models as representatives, we replace the audio with Gaussian white noise of the same length and input it into the audio-language models. As shown in Figure 4, the accuracy of the models dropped significantly when they received white noise, approaching random guessing. This indicates that the models received audio information. However, the results also show that even with noisy audio, the models were still able to derive some answers from the textual questions.
Error Analysis: We classify incorrect answers of four models using Gemini-2.5-flash and find that reasoning errors were the main source of error for most models. For example, the rate of reasoning errors of GPT-4o Audio was 86.68%, while the other three types of errors were relatively low.
For LALMs, there are various reasons that lead to model errors. Firstly, LALMs tend to answer the content of options rather than just A/B/C/D like text reasoning models, which results in different models having different preferences in answering questions. For example, Omni-R1 often adds line breaks, and some models cannot fully restate the options. Secondly, LALMs generally have randomness and instability, and different prompts and runs can also lead to significant differences in accuracy. Detailed analysis and case studies are provided in Appendix H and Appendix I. Finally, hallucination problem that LALMs encounter mainly manifests as the creation of non-existent sound events or options. Most of the problems stem from the insufficient granularity of model perception, leading to the omission of key content.
[…]
@@ 5 Conclusion
We introduce CMDAR, a Chinese benchmark for evaluating models on challenging, multi-scene, dynamically evolving audio reasoning tasks. CMDAR comprises 3,000 carefully curated question–answer pairs linked to diverse audio clips, covering five categories of complex reasoning across three question types. We evaluated 26 audio models, revealing that CMDAR poses substantial challenges across question formats and evaluation methods. Furthermore, we analyze gaps between open-source and closed-source models, and overall limitations in model audio understanding. Furthermore, we analyze the gap between open-source and closed-source models, the perceptual and reasoning capabilities of cascading caption models, and limitations of model audio reasoning, providing suggestions for future development.
Limitations
Our benchmark may has the following limitations.
Data. Our benchmark is based on monolingual Chinese data, which may have limitations due to language diversity. We hope that the audio community will focus on the development of Chinese audio reasoning capabilities.
Models. We have only tested a subset of LALMs. For the latest models, we will update the code leaderboard in a timely manner.
Evaluation. Although we conducted numerous experiments to validate the rationale for using LLM for assessment, it is undeniable that updates to the LLM framework can impact experimental results. We will update the leaderboard promptly and consider incorporating more robust metrics for a comprehensive evaluation of reasoning ability.
Analysis. Our analysis of the limitations and future development of LALMs is based on results from CMDAR and related work.
Ethics Statement
This study follows the relevant ethical guidelines of the institution and country, has obtained informed consent from all participants, and strictly protects their privacy and data security. The manual annotations in this work are solely for evaluation purposes in the ’Human–Large Language Model Protocol’ research, and the generated annotations will not be included in the published resources. All manual evaluations are conducted by expert reviewers selected from the research team according to a predefined assessment plan. No personal data have been collected. The research process has no conflicts of interest, and the results are accurate and reliable. Therefore, we believe this work does not raise significant ethical concerns.
Appendix A Task
A.1 Task Introduction
@@
A.2 Task Difficulty
Based on the experimental data in Table 1, Table 2, and Table 7, it can be found that: among the core task types of single-choice questions and open-ended questions, the scene understanding task is the most difficult, while the anomaly detection and event reasoning tasks also present certain challenges. Further analysis by question type shows that: in the context of open-ended questions, temporal reasoning is the most difficult task type, followed by the anomaly detection task in terms of difficulty; in the context of multiple-choice questions, the difficulty coefficient of the continuous scene reasoning task reaches the highest level. This result indicates that the current model still has significant limitations in its reasoning ability when dealing with dynamically switching continuous scenes, which is a key direction that requires focused breakthroughs in subsequent optimizations.
Appendix B Data Source
The CMDAR Metadata Movie Collection includes 500 carefully selected Chinese films, covering a wide range of genres and themes. It is designed to construct complex and dynamic scenarios, so as to meet the requirements of multi-threaded narratives and high production standards. Below is a detailed statistical breakdown of the collection. The 500 films in the CMDAR Metadata Movie Collection are categorized by common film genres and themes, with the statistical data presented in the Table 5.
Figure 5: Distribution of all types of tasks included in the three benchmark tests illustrates the diversity and complexity of CMDAR tasks.
@@ Appendix B Data Source
Table 6: Statistical information of the three benchmark tests indicates the complexity of the benchmarks through the longer duration of the audio and the rich semantic information.
Statistics Number
Total Questions 1500
Task 5
CMDAR-main Sub task 9
Average question length 30.68
Average option length 18.69
Average audio length 25.11
Total Questions 500
Task 5
CMDAR-open Average question length 41.47
Average reference length 175.53
Average audio length 25.14
Total Questions 1000
Task 3
Average question length 32.80
CMDAR-multi Average answer length 47.19
Average audio length 69.90
Average answer number 2.82
The CMDAR Metadata Movie Collection has the following characteristics:
@@
This data provides a solid foundation for the construction of the CMDAR benchmark and also offers abundant resources for relevant research.
Appendix C Prompt of Evaluation
"[Detailed Audio Description][Audio]"
"[Question][Question]"
"[The Start of Assistant 1s Answer][Assistant1]"
"[The End of Assistant 1s Answer]"
"[The Start of Assistant 2s Answer][Assistant2]"
"[The End of Assistant 2s Answer]"
"[System]"
"The two scores are separated by a space."
Assistant: {response}
"accuracy": "The accuracy of the response, whether it is factual"
Assistant:
{response}
@@ Appendix D Prompt of Data construction pipeline
Prompt of Data Generation in CMDAR-main System: You are a helpful AI assistant, designed to provide useful QA pair to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the provided video, audio, audio translation, audio description, and video description, create three questions about the audio. These should be three audio reasoning questions with options and answers. The audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. The questions must meet the following criteria: 1.They should be reasoning questions related to the audio. 2.The category of the questions should fall into one of the following: temporal reasoning task, social intent reasoning task, anomaly detection and safety task, and specify which category it belongs to. 3.The posed questions should include a thought chain explanation. 4.The questions cannot be answered solely by the text of the title, they must require reasoning based on both the question and the audio. 5.The questions should introduce some additional reasoning information that does not exist in the audio to increase difficulty. 6.Various information from the audio should not appear directly in the question text; just pose the question. 7.The answers provided should follow this format: Question:….. Question Category:….. Question Options:…… Question Answer:…… Assistant: {response}
@@
Prompt of Data generation in CMDAR-multi System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Based on the two provided relevant audio segments along with their respective translations, descriptions, and video descriptions, formulate multiple-choice audio reasoning questions-options-answers-reasoning chains for the audio pair. Audio translation 1 is [audio translation 1], audio description 1 is [audio description 1], video description 1 is [video description 1], audio translation 2 is [audio translation 2], audio description 2 is [audio description 2], video description 2 is [video description 2]. The questions must meet the following criteria: 1. The question must relate to audio pairs and cannot include any information or terms related to the video. 2. The category of the question should belong to one of the following categories, and the category must be specified: continuous scene reasoning task, multi-character interaction reasoning task, plot development reasoning task, anomaly detection and security task, time reasoning task. 3. The question should not have an answer that can be derived solely from the text title; it should require reasoning that combines the question and the audio. 4. The question should involve multi-step reasoning and be challenging, with options of moderate length for the answers. 5. The multiple-choice answers can independently reason for each audio’s description but must include at least one reasoning question that is related to both audio segments simultaneously. 6. Your provided answers must conform to the following format: Question:….. Question Category: …… Question Options: …… Question Answer: …… Question Reasoning Chain: …… Assistant: {response}
@@
Prompt of Data generation in CMDAR-open System: You are a helpful AI assistant, designed to provide useful answer to customers. The user will provide you with a video, audio, audio translation, audio description, and video description. Please refer to this desired output and provide your own unique response. User: Propose an open-ended question for audio based on the given video, audio, audio translation, audio description, and video description. The requirement is an open-ended reference answer thought chain pair, where the audio translation is [audio translation], the audio description is [audio description], and the video description is [video description]. At the same time, the questions raised must meet the following requirements: 1. The requirements are only open-ended reasoning questions related to audio, and the video and provided text are only for reference in question setting and cannot be used for question setting (already questions, answers) and thought chain analysis. The questions can only be answered by understanding and reasoning about the audio. 2. The category of the question belongs to one of the following categories, and indicate which task it belongs to: time reasoning task, scene understanding task, character relationship and social reasoning task, event reasoning task, anomaly detection and security task. 3. The answer to the question cannot be found solely from the text question, it needs to be combined with the question and audio for reasoning. 4. The answer to an open-ended question is not unique, just provide a reference answer. The answer you provide should meet the following format: Question: …. Question Category: …. Reference answer: …. Assistant: {response}